# Trustworthy Semantic-Enabled 6G Communication: A Task-oriented and Privacy-preserving Perspective

### v1

@@ V-C Transfer Learning for Task, Data and Channel Adaption
In learning-based privacy-preserving ToSC, the system needs to be retrained as either the task, data or channel varies. Transfer learning can be instrumental for task/data/channel adaptation, allowing communication systems to efficiently adapt to new domains using pre-existing knowledge. Task adaptation focuses on applying learned models to new but related tasks. In ToSC, this could involve using a model trained for one semantic communication task (e.g., image recognition) and adapting it for another (e.g., image retrieval). Transfer learning enables ToSC systems to quickly adjust to new tasks without the need for extensive retraining, thereby saving time and computational resources. Data adaptation is essential due to the variability of data types and sources. Transfer learning allows for the utilization of pre-trained models on one type of data (like text) and adapts them for different types (such as images or sensors data). This flexibility is particularly beneficial in multi-modal communication scenarios where different types of data need to be processed and transmitted seamlessly. Transfer learning can be also employed to adapt ToSC systems to varying channel conditions such as interference, signal attenuation, and mobility. By learning from data transmitted under different channel conditions, a ToSC system can predict and adjust its parameters for optimal performance, even in less-than-ideal transmission environments.
V-D Integrating Physical Layer Security
In ToSC, integrating physical layer security is paramount for preserving privacy by safeguarding the confidentiality, integrity, and availability of transmitted semantic information, especially when conventional security measures at higher layers are ineffective. Physical layer security strengthens communication systems by providing an additional defense layer against eavesdropping and jamming attacks. The resilience of the communication channel against compromise by adversaries could be enhanced by leveraging inherent properties of the communication medium, such as signal attenuation and channel randomness. However, there are still multifaceted challenges. Implementation often introduces overhead in bandwidth utilization, and energy consumption, necessitating careful consideration of trade-offs between security and system performance. Moreover, ToSC systems operate in dynamic and unpredictable environments, characterized by rapid fluctuations in channel conditions. Consequently, it is imperative to devise mechanisms resilient to adversarial threats and adaptable to diverse operational conditions, including variations in channel quality and eavesdropper locations.
@@ VI Conclusion
This article highlights the significance of transmitting task-specific essential information efficiently while addressing the critical issue of privacy preservation. It includes a comprehensive analysis of privacy-preserving strategies for ToSC, comparing feature perturbation methods like differential privacy and encryption with intrinsic security incorporation approaches such as adversarial learning and LBVQ. The research also explores experimental evaluations of these methods, assessing their performance and privacy protection capabilities. Finally, potential avenues for future study on privacy security and information trustworthiness are prospected.

### v2

@@ V-C Transfer Learning for Task, Data and Channel Adaption
In JSCC-based ToSC, the system needs to be retrained as either the task, data or channel varies. Transfer learning can be instrumental for task/data/channel adaptation, allowing communication systems to efficiently adapt to new domains using pre-existing knowledge. Task adaptation focuses on applying learned models to new but related tasks. Transfer learning enables ToSC systems to quickly adjust to new tasks without the need for extensive retraining, thereby saving time and computational resources. Data adaptation is essential due to the variability of data types and sources. Transfer learning allows for the utilization of pre-trained models on one type of data (like text) and adapts them for different types (such as images or sensors data). This flexibility is particularly beneficial in multi-modal communication scenarios where different types of data need to be processed and transmitted seamlessly. Transfer learning can be also employed to adapt ToSC systems to varying channel conditions such as interference, signal attenuation, and mobility. By learning from data transmitted under different channel conditions, a ToSC system can predict and adjust its parameters for optimal performance, even in less-than-ideal transmission environments [14].
@@ VI Conclusion
This paper highlights the significance of transmitting task-specific essential information efficiently while addressing the privacy preservation issue. It includes a comprehensive analysis of privacy-preserving strategies for ToSC, comparing feature perturbation methods like DP and encryption with intrinsic security incorporation approaches such as adversarial learning and LBVQ. Our research also explores experimental evaluations of these methods, assessing their performance and privacy protection capabilities. Finally, potential avenues for future study are provided.
Acknowledgments
The work is supported in part by the National Natural Science Foundation of China under Grant 62171262 and 62301328; in part by Shandong Provincial Natural Science Foundation under Grant ZR2021YQ47; in part by the Taishan Young Scholar under Grant tsqn201909043; in part by Major Scientific and Technological Innovation Project of Shandong Province under Grant 2020CXGC010109.